# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v3

@@ 5.6.1 Top Trading Cycles: A Variant for House Allocation with Existing Tenants
Abdulkadiroğlu and Sönmez (1999) introduces a second algorithm that generates the same outcome as YRMH–IGYT. Using a technical trick that considers individuals and houses as separate entities, this algorithm converts each chain of individuals in YRMH–IGYT into a cycle of individuals and houses, facilitating all trades through cycles. Hence, we call this algorithm Top Trading Cycles (TTC).
In this alternative procedure, not only individuals but also houses are represented as nodes in a directed graph. At any step, a directed link is formed from each individual to their most preferred house still available. Existing tenants have private ownership of their earmarked units, which is reflected by a directed link from each occupied house to its “owner”—the current occupant. For vacant units and those vacated by previously processed tenants, the given priority order determines property rights, represented by a directed link from each such house to the highest priority individual remaining in the procedure. This mapping allows us to adopt the same dynamics of the G–TTC algorithm for house allocation with existing tenants.
While YRMH–IGYT and TTC generate the same outcome for any given problem, and thereby induce the same direct mechanism, they offer distinct advantages. As presented in Section 5.6, YRMH–IGYT is a minimalist refinement of the SSD with Squatting Rights, making it an easier sell for institutions using this mechanism. In contrast, TTC, though more abstract, not only has a closer connection with earlier literature but also readily accommodates unit-specific priority queues for vacant units. As later discussed in Section 6.8, this observation forms the basis of the TTC variant we introduced in Abdulkadiroğlu and Sönmez (2003) for school choice, as part of one of the “discovery-invention cycles” highlighted in this monograph.
@@ 5.6.2 Contrast with a Technocratic Design Based on the Core
In the late 1990s, before it was published, I presented Abdulkadiroğlu and Sönmez (1999) at several universities. Many economic theorists in the audiences inquired why we felt compelled to design a new mechanism for the house allocation with existing tenants model when we could adopt a simpler “technocratic” solution based on the G–TTC algorithm:
First, convert the problem into a housing market by randomly assigning vacant houses to newcomers as their initial endowments using a uniform distribution. Since existing tenants already have their initial endowments, this step generates a housing market as long as there are at least as many newcomers as vacant houses. [^42] If there are more newcomers than vacant houses, some newcomers will not be assigned initial endowments and will remain unassigned under the proposed mechanism. The outcome of this technocratic solution is the unique core allocation of the resulting housing market. For lack of a better term, let’s refer to this mechanism as the core-based technocratic mechanism. [^43] When there are the same number of houses as individuals and all individuals are newcomers, this mechanism is called the core from random endowments (CRE). Interestingly, for any house allocation problem, the CRE generates the same probability distribution over matchings as the RSD (Knuth, 1996; Abdulkadiroğlu and Sönmez, 1998).
By construction, the core-based technocratic mechanism satisfies individual rationality. Moreover, by directly invoking results from earlier housing market literature, we see that it also satisfies ex-post Pareto efficiency and strategy-proofness. So, what was the point of our scholarly inquiry into the on-campus housing problem? Earlier basic theory already provided a solution to this practical application.
@@
This answer was not bad, and at the time, it seemed to satisfy the audience. However, deep inside, I was not fully satisfied with my response. Something was amiss. Nevertheless, the paper was accepted for publication, and my interest shifted to other projects that seemed more important at the time. It would take a few years and starting what became a very fruitful collaboration with another young economist, Utku Ünver, to return to this question.
@@
For any house allocation problem with existing tenants, the outcome of the core-based technocratic mechanism is same as the outcome of the YRMH–IGYT mechanism with any sequence of individuals that (i) prioritizes each newcomer over each existing tenant, and (ii) randomly prioritizes newcomers among themselves with uniform distribution. [^44] When each newcomer is prioritized over each existing tenant, the relative processing sequence of existing tenants among themselves becomes immaterial under the YRMH–IGYT (or TTC) algorithm.
@@
While the mechanism’s description doesn’t immediately reveal this bias, its source becomes clear upon reflection. To transform the problem into a housing market and leverage its well-understood solution, the core, vacant houses are exclusively allocated to newcomers during the artificial construction of an initial endowment. This technical construct deprives existing tenants of potential trade opportunities, including direct consumption, associated with these units. Despite the collective ownership of the vacant houses by all individuals, existing tenants effectively lose these property rights under the core-based technocratic mechanism, being confined to direct consumption or trades involving their earmarked units. While this approach may be desirable under some policies, it clearly isn’t in others.
Perhaps the most concerning aspect is that this bias is subtly embedded in the system, compromising transparency and creating opportunities for exploitation by ill-intentioned decision makers or for unintended consequences by well-meaning ones (see Section 13 for further discussion on these behavioral considerations). This phenomenon is not uncommon in applications where a solution to a complex problem is derived from a well-understood solution to a simpler case. With its emphasis on custom-made theory tailored to specific applications, minimalist market design is more resilient against such failures.
@@ 6 School Choice
Although Abdulkadiroğlu and Sönmez (1999) was motivated by campus housing at various universities, I never pursued policy initiatives with their housing offices to reform flawed mechanisms like SSD (or RSD) with Squatting Rights. As the project progressed, my interest shifted more towards the theory of these problems, largely due to the richness and elegance of the underlying analytical framework. By the time we developed YRMH–IGYT, another application—Turkish college admissions—seemed far more important from a policy perspective.
The evolution of minimalist market design in actual policy can be traced through a series of holistic research and policy initiatives concerning the allocation of college seats in Turkey (Balinski and Sönmez, 1999) and K-12 public school seats in the U.S. (Abdulkadiroğlu and Sönmez, 2003; Chen and Sönmez, 2006; Ergin and Sönmez, 2006) from the late 1990s to mid-2000s. Based on these experiences, along with those from another early application—kidney exchange, presented in Section 7—I began deliberately and rigorously applying minimalist market design in my research and policy initiatives starting in the early 2010s.
@@ 6.1 Student Placement in Turkey
In Turkey, the student placement problem for college admissions takes a more complex form. Although all college seats are assigned through a centralized mechanism using a single standardized test, the system constructs multiple field-specific rankings by assigning different weights to sections of the test. Each college is exogenously mapped to one of these rankings based on its field. For instance, the ranking for engineering schools places a higher weight on mathematics questions than the one for medical schools. As a result, SSD is no longer suitable for solving the problem, as there is no single priority ranking.
@@ 6.2 No Justified Envy
An allocation of school seats to students satisfies no justified envy, [^45] This axiom is called fainess in Balinski and Sönmez (1999). Referral to this axiom through an absence of justified envy is first done in Abdulkadiroğlu and Sönmez (2003), where it is called elimination of justified envy. if, there is no school $s$ and two distinct students $i$, $j$ such that,
@@
    student $i$ has a higher score than student $j$ for the standardized test for school $s$.
@@
Given a student placement problem, a vector of cutoff scores supports an allocation of school seats to students if,
@@
    Any student $i$ who is assigned a seat at school $s$ has a score at least as high as the cutoff score for school $s$, and
@@
    any student $i$ who has a score at least as high as the cutoff score of any given school $s$ weakly prefers their assignment over school $s$.
@@
Therefore, the practice of announcing outcomes based on cutoff scores—a practice common worldwide—is a feature of mechanisms that satisfy no justified envy.
@@
When the priority ranking of students is identical at all institutions, SSD is the only direct mechanism that satisfies no justified envy and Pareto efficiency.
@@
$\begin{array}[]{lll}\mbox{Alp}:&X-Y&X:\;\;\mbox{Banu, Cora, Alp}\\ \mbox{Banu}:&Y-X&Y:\;\;\mbox{Alp, Banu, Cora}\\ \mbox{Cora}:&X-Y&\end{array}$
Since Alp and Banu are each the highest-priority student at their respective second-choice colleges, neither can receive a less preferred assignment under any allocation that satisfies no justified envy. Thus, the two seats at $X$ and $Y$ must be shared between Alp and Banu, leaving Cora unassigned.
However, Cora’s presence is not without consequence for Alp and Banu. Since she is ranked higher than Alp at college $X$, Alp cannot be assigned a seat at his first choice under any allocation that satisfies no justified envy. As a result, Alp must receive a seat at his second choice, which means Banu must also receive a seat at her second-choice college, $Y$. But this allocation is Pareto dominated by the allocation where Alp and Banu receive their first choices, establishing the potential incompatibility between no justified envy and Pareto efficiency.
@@ 6.3 Multi-Category Serial Dictatorship
For some reason, policymakers often choose SSD as the starting point for their allocation mechanisms, even in settings where it cannot be directly applied. Similar to many institutions that use a modification of SSD—specifically, SSD with Squatting Rights—for practical applications in house allocation with existing tenants (see Section 5.5, Turkish authorities have adopted an iterative procedure that incorporates SSD as its core engine to address the more complex issue of college admissions with field-specific student rankings.
@@
This process is repeated with the modified preferences until no student is assigned seats at multiple colleges, at which point the tentative assignments are finalized. Since the preference ranking of any student holding seats at multiple colleges strictly shortens with each step, the procedure terminates in a finite number of iterations. Balinski and Sönmez (1999) refers to the resulting direct mechanism as the multi-category serial dictatorship (MCSD).
Let’s illustrate this mechanism with an example.
@@
There are five students: Alp, Banu, Cora, Diya, and Ezra. There is an engineering college, $X$, with two seats and two medical colleges, $Y$ and $Z$, each with one seat. Students are evaluated in two categories: math and science. They are priority-ranked by their math scores at college $X$ and by their science scores at colleges $Y$ and $Z$. The preferences of the students over their acceptable colleges and their priority rankings for math and science are given as follows:
$\begin{array}[]{llllll}\mbox{Alp}:&Y-X&\mbox{Diya}:&X-Y&\mbox{Math}:&\;\mbox{Alp, Banu, Cora, Diya, Ezra}\\ \mbox{Banu}:&X-Y-Z&\mbox{Ezra}:&Y-Z-X&\mbox{Science}:&\;\mbox{Alp, Diya, Cora,Banu, Ezra}\\ \mbox{Cora}:&X-Z-Y&&&&\end{array}$
@@
Let’s refer to the SSD induced by the math ranking as SSD-math, and the SSD induced by the science ranking as SSD-science. Throughout the procedure, the two seats at $X$ are allocated using SSD-math, while the two seats at $Y$ and $Z$ are allocated using SSD-science.
@@
In Step 1, SSD-math assigns the two seats at $X$ to Alp and Banu, the two students with the highest math ranking. SSD-science assigns the seat at $Y$ to the highest science-ranking student, Alp, and the seat at $Z$ to the third science-ranking student, Cora, since it is not acceptable to the second science-ranking student, Diya. Since Alp holds seats at two colleges, the procedure needs to continue with Step 2. Holding at least one seat each, the preferences of Alp, Banu, and Cora are truncated after their most preferred tentative assignments. This results in the following modified list of preferences for Step 2:
$\begin{array}[]{llll}\mbox{Alp}:&Y&\mbox{Diya}:&X-Y\\ \mbox{Banu}:&X&\mbox{Ezra}:&Y-Z-X\\ \mbox{Cora}:&X-Z&&\end{array}$
@@
In Step 2, SSD-math assigns the two seats at $X$ to Banu and Cora, the second and third highest in math ranking, since school $X$ is no longer acceptable to the highest-ranking student, Alp. Just as in Step 1, SSD-science assigns the seat at $Y$ to the highest science-ranking student, Alp, and the seat at $Z$ to the third science-ranking student, Cora, since it is not acceptable to the second science-ranking student, Diya. Since Cora holds seats at two colleges, the procedure needs to continue with Step 3. Holding at least one seat each, the preferences of Alp, Banu, and Cora are truncated after their most preferred tentative assignments. With only a single acceptable choice remaining for each, this modification does not further truncate the preferences of Alp and Banu, but only Cora’s, resulting in the following further modified list of preferences for Step 3:
$\begin{array}[]{llll}\mbox{Alp}:&Y&\mbox{Diya}:&X-Y\\ \mbox{Banu}:&X&\mbox{Ezra}:&Y-Z-X\\ \mbox{Cora}:&X&&\end{array}$
@@
In Step 3, SSD-math assigns the two seats at $X$ to Banu and Cora, just as in Step 2. SSD-science assigns the seat at $Y$ to the highest science-ranking student, Alp, and the seat at $Z$ to the fifth science-ranking student, Ezra, since it is not acceptable to the second, third, and fourth science-ranking students, Diya, Cora, and Banu. Since no one holds seats at multiple colleges, the procedure terminates, finalizing the latest tentative assignments.
@@
When SSD is tweaked into SSD with Squatting Rights for house allocation with existing tenants, as presented in Section 5.5, the modification feels somewhat “forced,” making the design’s shortcomings rather evident. In contrast, MCSD—though also an SSD-based design—appears quite compelling at first glance. For one, it satisfies no justified envy, which is the most critical axiom in the context of Turkish college admissions. However, the following simple example reveals that constructing a mechanism for this more complex version of student placement based on SSD also incurs an efficiency cost.
@@
$\begin{array}[]{llll}\mbox{Alp}:&X-Y&\mbox{Math}:&\mbox{Banu, Alp}\\ \mbox{Banu}:&Y-X&\mbox{Science}:&\mbox{Alp, Banu}\\ \end{array}$
@@
In Example 3, we observed that some problems have no outcome satisfying both no justified envy and Pareto efficiency. However, this is not the case in Example 5. Thus, the potential welfare loss under MCSD cannot be attributed to this fundamental incompatibility between the two axioms. Although not immediately evident, the root cause of MCSD’s failures lies in its reliance on SSD as its core mechanism. To understand why, it is helpful to relate the student placement model to the celebrated college admissions model by Gale and Shapley (1962).

### v4

@@ 5.6.1 Top Trading Cycles: A Variant for House Allocation with Existing Tenants
Abdulkadiroğlu and Sönmez (1999) introduces a second algorithm that yields the same outcome as YRMH–IGYT. Employing a technical construct that treats individuals and houses as separate entities, this algorithm transforms each chain of individuals in YRMH–IGYT into a cycle of individuals and houses, thereby implementing all trades through cycles. Accordingly, we denote this algorithm as the House Allocation variant of Top Trading Cycles (HA–TTC). [^51] The term HA–TTC mechanism is used interchangeably with the YRMH–IGYT mechanism in this monograph.
In this alternative procedure, not only individuals but also houses are represented as nodes in a directed graph. At any step, a directed link is formed from each individual to their most preferred house still available. Existing tenants have private ownership of their earmarked units, which is reflected by a directed link from each occupied house to its “owner”—the current occupant. For vacant units and those vacated by previously processed existing tenants, the given priority order determines property rights, represented by a directed link from each such house to the highest priority individual remaining in the procedure. This mapping enables dynamics analogous to those of the G–TTC algorithm for house allocation with existing tenants.
As we will see in several other cases—and discuss in depth in Section 13.4—different algorithms that induce the same mechanism, or even different framings of a mechanism (and, more broadly, of an institution), may offer distinct advantages for practical design. In this vein, while YRMH–IGYT and HA–TTC generate the same outcome for any given problem and thus induce the same direct mechanism, they offer different benefits. As presented in Section 5.6, YRMH–IGYT is a minimalist refinement of SSD with Squatting Rights, making it easier to adopt for institutions already using that mechanism. By contrast, HA–TTC—though more abstract—connects more directly to earlier literature and readily accommodates unit-specific priority queues for vacant units, a feature absent in YRMH–IGYT. Moreover, as discussed in Section 6.8, this added flexibility underlies the school choice variant of the Top Trading Cycles algorithm introduced in Abdulkadiroğlu and Sönmez (2003), as part of one of the “discovery—invention cycles” highlighted in this monograph (see Figure 36).
@@ 5.6.2 Contrast with a Technocratic Design Based on the Core
In the late 1990s, before its publication, I presented Abdulkadiroğlu and Sönmez (1999) at several universities. Many economic theorists in the audience asked why we felt compelled to design a new mechanism for house allocation with existing tenants when a simpler “technocratic” solution based on the G–TTC algorithm could be adopted:
First, convert the problem into a housing market by randomly assigning vacant houses to newcomers as their initial endowments using a uniform distribution. Since existing tenants already have their initial endowments, this step yields a housing market as long as the number of newcomers is at least as large as the number of vacant houses. [^52] If there are more newcomers than vacant houses, some newcomers will not be assigned initial endowments and will remain unassigned under the proposed mechanism. Next, select the unique core allocation of the resulting housing market as the outcome. For lack of a better term, I refer to this mechanism as the core-based technocratic mechanism. [^53] When the number of houses equals the number of individuals and all individuals are newcomers, this mechanism is called the core from random endowments (CRE). Interestingly, for any house allocation problem, the CRE generates the same probability distribution over matchings as RSD (Knuth, 1996; Abdulkadiroğlu and Sönmez, 1998).
By construction, the core-based technocratic mechanism satisfies individual rationality. Moreover, by directly invoking results from earlier housing market literature, the mechanism also satisfies ex-post Pareto efficiency and strategy-proofness. What, then, was the point of our scholarly inquiry into the on-campus housing problem? After all, basic theory had already provided a solution to this practical application.
@@
This answer was not bad, and at the time, it seemed to satisfy the audience. However, deep inside, I was not fully satisfied with my response. Something was amiss. Nevertheless, the paper was accepted for publication, and my interest shifted to other projects that seemed more important at the time. It would take a few years, and embarking on what would become a very fruitful collaboration with another young economist, Utku Ünver, before I returned to this question.
@@
For any house allocation problem with existing tenants, the outcome of the core-based technocratic mechanism is same as the outcome of the YRMH–IGYT mechanism with any sequence of individuals that (i) prioritizes each newcomer over each existing tenant, and (ii) randomly prioritizes newcomers among themselves with uniform distribution. [^54] When each newcomer is prioritized over each existing tenant, the relative processing sequence of existing tenants among themselves becomes immaterial under the YRMH–IGYT (or HA–TTC) algorithm.
@@
While the mechanism’s description does not immediately reveal this bias, its source becomes clear upon reflection. To transform the problem into a housing market and leverage its well-understood solution—the core—vacant houses are assigned exclusively to newcomers during the artificial construction of an initial endowment. This technical construct deprives existing tenants of potential trade opportunities, including direct consumption, associated with these units. Even when existing tenants may have higher or equal claims to vacant houses, as is common for upperclassmen in on-campus housing, they effectively lose these property rights under the core-based technocratic mechanism, being confined to direct consumption or trades involving only their earmarked units.
Perhaps the most concerning aspect is that this bias is subtly embedded in the system, compromising transparency and creating opportunities for exploitation by ill-intentioned decision-makers or for unintended consequences by well-meaning ones (see Section 13 for further discussion on these behavioral considerations). This phenomenon is not uncommon in applications where a solution to a complex problem is derived from a well-understood solution to a simpler case. With its emphasis on custom-made theory tailored to specific applications, minimalist market design is more resilient against such failures.
@@ 6 School Choice
Although Abdulkadiroğlu and Sönmez (1999) drew its motivation from campus housing at various universities, I never pursued policy initiatives with their housing offices to reform flawed mechanisms like SSD (or RSD) with Squatting Rights. As the project progressed, my interest shifted more towards the theory of these problems, largely due to the richness and elegance of the underlying analytical framework. By the time we developed YRMH–IGYT, another application—Turkish college admissions—seemed far more important from a policy perspective.
The evolution of minimalist market design in actual policy can be traced through a series of holistic research and policy initiatives concerning the allocation of college seats in Turkey (Balinski and Sönmez, 1999) and K–12 public school seats in the U.S. (Abdulkadiroğlu and Sönmez, 2003; Chen and Sönmez, 2006; Ergin and Sönmez, 2006) from the late 1990s to mid-2000s. Based on these experiences, along with those from another early application—kidney exchange, presented in Section 7—I began deliberately and rigorously applying minimalist market design in my research and policy initiatives starting in the early 2010s.
@@ 6.1 Student Placement in Turkey
In Turkey, the student placement problem for college admissions takes a more complex form. Although all college seats are assigned through a centralized mechanism based on a single standardized test, the system constructs multiple field-specific rankings by assigning different weights to sections of the test. Each college is exogenously mapped to one of these rankings according to its field. For example, the ranking for engineering schools places greater weight on mathematics questions than the ranking for medical schools. As a result, SSD is no longer suitable for this problem, since no single priority ranking applies to all schools.
@@ 6.2 No Justified Envy
An allocation of school seats to students satisfies no justified envy, [^55] This axiom is called fairness in Balinski and Sönmez (1999). In a simpler setting—where a single priority ranking applies at all institutions and the number of items equals the number of individuals—Svensson (1994) earlier introduced this axiom as weak fairness. A reference to this axiom via the absence of justified envy first appears in Abdulkadiroğlu and Sönmez (2003), where it is called elimination of justified envy. if there is no school $s$ and two distinct students $i$ and $j$ such that,
@@
    student $i$ has a higher score than student $j$ on the standardized test for school $s$.
@@
Given a student placement problem, a vector of cutoff scores supports an allocation of school seats to students if
@@
    every student $i$ who is assigned a seat at school $s$ has a score at least as high as the cutoff score for school $s$, and
@@
    every student $i$ whose score is at least as high as the cutoff score of some school $s$ weakly prefers their assignment to school $s$.
@@
Therefore, the practice of announcing outcomes based on cutoff scores—a practice common worldwide—is a hallmark of mechanisms that satisfy no justified envy.
@@
When the priority ranking of students is identical at all institutions, SSD is the unique direct mechanism that satisfies no justified envy and Pareto efficiency.
@@
[…]
Since Alp and Banu are each the highest-priority student at their respective second-choice colleges, neither can be assigned a less-preferred school under any allocation that satisfies no justified envy. Thus, the two seats at $X$ and $Y$ must be shared between Alp and Banu, leaving Cora unassigned.
However, Cora’s presence is not without consequence for Alp and Banu. Because she is ranked higher than Alp at college $X$, Alp cannot be assigned his first-choice college $X$ under any allocation that satisfies no justified envy. Consequently, Alp must receive a seat at his second-choice college $Y$, which in turn forces Banu to receive a seat at her second-choice college $X$. Yet this allocation is Pareto dominated by the one in which Alp and Banu receive their first choices, establishing the potential incompatibility between no justified envy and Pareto efficiency. (See Figure 7 for an illustration.)
Figure 7: Efficiency cost of no justified envy (NJE). Panel (a) presents the environment: individuals A, B, and C rank-order preferences over assignments X, Y, and $\varnothing$ (unmatched). Each school’s priority for a student is indicated alongside the student’s preferences; for example, under A, $X(3)$ indicates that A has third priority at $X$. Because both schools are acceptable to all, NJE requires that no student be assigned an outcome worse than a school that ranks them first. Thus, A and B cannot be placed below their second choices; consequently, $X$ and $Y$ must be allocated between them (Panel (b)), leaving C unmatched (Panel (c)). Since C prefers $X$ to its assignment $\varnothing$ and outranks A at $X$, NJE rules out assigning A to $X$ (Panel (d)); therefore, A receives $Y$ (Panel (e)), which in turn implies that B receives $X$ as the remaining option (Panel (f)). This allocation gives A and B their second choices and keeps C unmatched; it is Pareto dominated by the allocation in which A and B obtain their first choices while C remains unmatched.
@@ 6.3 Multi-Category Serial Dictatorship
What the mechanism SSD represents is so deeply intertwined with the use of priority rankings in unit-demand assignment problems with indivisible goods that policymakers often take SSD as the natural starting point for allocation procedures, even in settings where it cannot be directly applied. Much like many institutions that use a modification of SSD—specifically, SSD with Squatting Rights—for practical applications in house allocation with existing tenants (see Section 5.5), Turkish authorities adopted an iterative procedure that embeds SSD as its core engine to address the more complex problem of college admissions with field-specific student rankings.
@@
This process is repeated with the modified preferences until no student holds seats at multiple colleges, at which point the tentative assignments are finalized. Because the preference list of any student with multiple seats strictly shortens at each step, the procedure terminates after a finite number of iterations. The resulting direct mechanism is known as the multi-category serial dictatorship (MCSD).
We now illustrate this mechanism with an example.
@@
There are five students: Alp, Banu, Cora, Diya, and Ezra. There is an engineering college, $X$, with two seats, and two medical colleges, $Y$ and $Z$, each with one seat. Students are evaluated in two categories: math and science. They are priority-ranked by their math scores at college $X$ and by their science scores at colleges $Y$ and $Z$. The preferences of the students over their acceptable colleges and their priority rankings in math and science are given as follows:
[…]
@@
We refer to the SSD induced by the math ranking as SSD-math, and the SSD induced by the science ranking as SSD-science. Throughout the procedure, the two seats at $X$ are allocated using SSD-math, while the seats at $Y$ and $Z$ are allocated using SSD-science.
@@
In Step 1, SSD-math assigns the two seats at $X$ to Alp and Banu, the students with the two highest math rankings. SSD-science assigns the seat at $Y$ to Alp, the highest-ranked in science, and the seat at $Z$ to Cora, the third-ranked in science, since it is not acceptable to Diya, who is ranked second. Because Alp holds seats at two colleges, the procedure continues to Step 2. For students holding at least one seat, their preferences are truncated after their most preferred tentative assignment, giving the following modified list of preferences:
[…]
@@
In Step 2, SSD-math assigns the two seats at $X$ to Banu and Cora, the second- and third-ranked in math, since $X$ is no longer acceptable to Alp. As in Step 1, SSD-science assigns the seat at $Y$ to Alp and the seat at $Z$ to Cora, since Diya, ranked second in science, does not find $Z$ acceptable. Because Cora now holds seats at two colleges, the procedure continues to Step 3. Again, for students holding at least one seat, preferences are truncated after their most preferred tentative assignment. This modification does not further affect Alp and Banu, but reduces Cora’s list, resulting in:
[…]
@@
In Step 3, SSD-math assigns the two seats at $X$ to Banu and Cora, just as in Step 2. SSD-science assigns the seat at $Y$ to Alp and the seat at $Z$ to Ezra, the fifth-ranked in science, since it is not acceptable to Diya, Cora, or Banu, ranked second through fourth. Because no student holds seats at multiple colleges, the procedure terminates, finalizing the tentative assignments. See Figure 8 for the execution of MCSD.
Figure 8: Execution of the MCSD mechanism in Example 5. The top panel illustrates a setting with five students (A, B, C, D, E) and three schools ($X$, $Y$, $Z$), where school $X$ belongs to the Math category and offers two seats, while schools $Y$ and $Z$ belong to the Science category and offer one seat each. In each category, students are priority ordered by their field-specific scores, with their school preferences ranked vertically beneath them. Within these preference lists, schools from different categories are shown in faint gray text. At each step, in both categories, seats are tentatively assigned using simple serial dictatorship (SSD), and any school ranked lower than a student’s best tentative assignment is removed from their preferences. The procedure terminates in Step 3, when no student holds more than one tentative assignment, finalizing the outcome.
@@
When SSD is modified into SSD with Squatting Rights for house allocation with existing tenants, as presented in Section 5.5, the adjustment feels somewhat “forced,” and the design’s shortcomings become rather evident. In contrast, MCSD—though also an SSD-based design—appears quite compelling at first glance. For one, it satisfies no justified envy, the most critical axiom in the context of Turkish college admissions. Yet the following simple example, depicted in Figure 9, shows that the reliance on SSD in the execution of MCSD also entails an efficiency cost.
@@
[…]
@@
Figure 9: MCSD is not second-best subject to no justified envy. The left panel represents the setting in Example 6, with two students (A and B) and two schools ($X$ and $Y$), each with one seat, where $X$ is in the Math category and $Y$ is in the Science category. In each category, students are priority-ordered by their field-specific scores, with their school preferences ranked vertically beneath them. Within these preference lists, schools from different categories are shown in faint gray. The right panel shows Step 1 of the execution of MCSD, in which seats in both categories are tentatively assigned via simple serial dictatorship. Because no student holds more than one tentative assignment, the procedure terminates immediately at Step 1, finalizing the tentative assignments. Under MCSD, both students receive their second choices. Another outcome that satisfies no justified envy—one that assigns both students their first choices—Pareto-dominates the outcome of MCSD.
In Example 4, we saw that some problems admit no outcome satisfying both no justified envy and Pareto efficiency. By contrast, this is not the case in Example 6. Hence, the potential welfare loss under MCSD cannot be attributed to this fundamental incompatibility between the two axioms. Although not immediately apparent, the root cause of MCSD’s shortcomings lies in its excessive reliance on SSD as a procedural subroutine. To see why, it is helpful to relate the student placement model to the celebrated college admissions model of Gale and Shapley (1962).